# Rotatable Antenna Aided Mixed Near-Field and Far-Field Communications in the Upper Mid-Band: Interference Analysis and Joint Optimization

### v2

@@ III-A Closed-form Expressions for Both Interference Terms
We first provide a closed-form approximation for the rotation-aware normalized near-field inter-user interference in (15), which is shown below.
@@
The rotation-aware normalized near-field inter-user interference $\rho_{\rm NN}(\phi,\theta_{k},r_{k},\theta_{i},r_{i})$ in (15) can be approximated as:
@@
Moreover, $\widehat{C}(\beta_{1},\beta_{2})={C}(\beta_{1}+\beta_{2})-{C}(\beta_{1}-\beta_{2})$ and $\widehat{S}(\beta_{1},\beta_{2})={S}(\beta_{1}+\beta_{2})-{S}(\beta_{1}-\beta_{2})$, where $C(\cdot)$ and $S(\cdot)$ are the Fresnel integrals, defined as $C(x)=\int^{x}_{0}\cos(\frac{\pi}{2}t^{2})\text{d}t$ and $S(x)=\int^{x}_{0}\sin(\frac{\pi}{2}t^{2})\text{d}t$.
@@
Next, we provide the closed-form approximation for the rotation-aware normalized mixed-field inter-user interference.
@@
The rotation-aware normalized mixed-field inter-user interference $\rho_{\rm NF}(\phi,\theta_{k},r_{k},\psi_{m})$ in (16) is approximated as
@@
It is worth highlighting that Theorem 1 generalizes Theorem 2, reducing to it as $r_{k}$ (or $r_{i}$) approaches infinity, i.e., $r_{k}(r_{i})\rightarrow\infty$ . Moreover, since both closed-form expressions accounting for the near-field and mixed-field interference are derived capitalizing on the Fresnel integrals, they share the same properties of function $G(\cdot)$, which facilitate the subsequent interference analysis, as detailed below.
@@
      For $\rho_{\rm NN}(0,\theta_{k},r_{k},\theta_{i},r_{i})=G(\bar{\beta}_{1},\bar{\beta}_{2})$ [13], where
@@
      For $\rho_{\rm NF}(0,\theta_{k},r_{k},\psi_{m})=G(\tilde{\beta}_{1},\tilde{\beta}_{2})$ [10], where
@@
Theorems 1 and 2 reveal an important insight: unlike the conventional fixed-antenna systems, the rotation angle $\phi$ introduces an additional DoF for controlling the values of $\beta_{1}$ and $\beta_{2}$, and consequently, the interference levels $G(\cdot)$. This manifests that the added flexibility enables the suppression of both near-field and mixed-field interference through strategic adjustment of the rotation angle $\phi$. In the subsequent subsections, we analytically demonstrate how the rotation angle $\phi$ affects the near-field and mixed-field interference, respectively, and hence the overall system performance.
@@ III-B Near-field Inter-user Interference
For two near-field users located at the same angle, i.e., $\theta_{k}=\theta_{i}=\theta$, but different distances, i.e., $r_{k}\neq r_{i}$, the rotation-aware normalized near-field interference in (17) depends solely on $\beta_{2}$ in (1), such that $\rho_{\rm NN}\approx G(0,\beta_{2})$. The effect of array rotation on this interference can be divided into the following two cases:
@@
    $\phi^{*}=\begin{cases}\min(\theta-\frac{\pi}{2},\phi_{\rm max}),&{\rm if}\leavevmode\nobreak\ \theta>\frac{\pi}{2},\\ \max(\theta-\frac{\pi}{2},\phi_{\rm min}),&{\rm if}\leavevmode\nobreak\ \theta<\frac{\pi}{2}.\end{cases}$ (24)
@@
To demonstrate the benefits of RAs in mitigating the near-field inter-user interference, we first validate the accuracy of the proposed approximation in Theorem 1 and the results derived in Lemma 1. This validation is presented in Fig. 5, where the Fresnel-based approximation is shown to closely match the exact normalized interference given in (15). Moreover, the derived rotation angles align closely with the numerical values across various spatial-angle settings. Next, we investigate the effects of the rotation angle on the achievable sum-rate with respect to the spatial angle and distance of the near-field user. This analysis is detailed in Figs. 55, where the system parameters are set to $N=129$, $f=24$ GHz, and $K=2$ (i.e., near-field users NU1 and NU2). We see that RA-enabled systems consistently outperform their fixed-antenna counterparts across all tested angles and distances. We explore the effect of rotation on the achievable sum-rate from both angular and distance perspectives as follows:
@@
  • Effect of angle given the same distance: Fig. 5 illustrates that when two near-field users are located at the same distance, the performance gain from RAs depends heavily on their angular separation. Notably, significant gains are observed when the angular separation is small, where interference is inherently strong, and RAs can effectively reduce it by adjusting the array’s orientation. As the angular separation increases, both RA-enabled and fixed-antenna systems experience improved performance due to the attenuation of interference caused by greater angular diversity.
@@ III-C Mixed-field Inter-user Interference
When the far-field user and the near-field user are at the same angle, i.e., $\psi=\theta$, the rotation-aware normalized mixed-field interference in (19) depends only on $\beta_{2}$ in (2), such that $\rho_{\rm NF}\approx G(0,\beta_{2})$. The effect of rotation on this interference can be characterized as follows:
@@
    $\phi^{*}=\begin{cases}\min(\psi-\frac{\pi}{2},\phi_{\rm max}),&{\rm if}\leavevmode\nobreak\ \psi>\frac{\pi}{2},\\ \max(\psi-\frac{\pi}{2},\phi_{\rm min}),&{\rm if}\leavevmode\nobreak\ \psi<\frac{\pi}{2}.\end{cases}$ (27)
@@
  • When the near-field and far-field users share the same spatial angle but are not at the boresight direction (i.e., $\theta=\psi\neq\frac{\pi}{2}$), performance gains from array rotation depend on the spatial angle, which diminishes when the angle approaches the boresight direction from either side.
@@ IV-A Inner Problem: Power Allocation Optimization
where the gradient of $B_{1}$ in terms of $\{P_{{\rm N},k}\}$ is given by (32), as shown on the bottom of the next page. The resultant problem is convex with respect to $\{P_{{\rm N},k}\}$, and thus can be efficiently solved by standard convex solvers such as CVX.
@@
IV-B Outer Problem: Rotation Angles Optimization
@@
$f_{\rm fit}(\bm{\phi}^{(t)}_{s})=\sum_{k\in\mathcal{K}}\leavevmode\nobreak\ R_{k}(\bm{\phi}^{(t)}_{s})-\xi\mathbb{P}(\bm{\phi}^{(t)}_{s})$, (39)
@@ V-A Sum-rate Versus Maximum Transmit Power
In Fig. 9, we illustrate the effect of the transmit power on the sum-rate performance of various schemes. First, it is observed that the proposed scheme closely matches the performance of its combination with ZF digital beamforming and consistently outperforms the benchmark schemes, highlighting the effectiveness of RAs in enhancing the mixed-field communication performance. Second, compared to the RA+EPA scheme, which optimizes rotation alone, joint optimization of rotation and power allocation yields a higher sum-rate, underscoring the efficacy of the integrated approach. Third, subarray-wise rotation delivers significant performance gains over the array-wise rotation, owing to its superior interference mitigation capabilities enabled by greater design flexibility. Additionally, one interesting observation is that the FA+ZF scheme underperforms the proposed method across all transmit power levels. This is because, while FP+ZF effectively nullifies the near-field inter-user interference, it simultaneously suffers from the complex mixed-field interference caused by far-field users, resulting in notable performance degradation, especially when the mixed-field interference is pronounced.
@@ VI Conclusion
In this paper, we proposed leveraging RAs to mitigate near-field and mixed-field interference by exploiting the spatial DoF enabled by RA array rotation. For the special case where all subarrays share a uniform rotation angle, we derived the closed-form expressions for the rotation-aware near-field interference and the rotation-aware mixed-field interference using the Fresnel integrals. In particular, we confirmed that array rotation effectively suppresses complex near-field and mixed-field interference, thus achieving substantial performance improvements in mixed-field communication scenarios. For the general case with subarray-wise rotation, we have developed a double-layer algorithm, where the inner layer optimizes power allocation via SCA, and the outer layer determines subarray rotation angles via PSO. Numerical results demonstrated that the RA-enabled system substantially outperforms conventional fixed-antenna systems and underscore the superior effectiveness of our proposed joint design compared to benchmark schemes. In the future, it would be of great interest to investigate hardware-efficient hybrid beamforming design tailored for RAs and novel mixed-field interference suppression techniques to further enhance the performance of mixed-field communication systems.
Appendix A Proof of Theorem 1
@@ Appendix B Proof of Lemma 1
Based on the properties of function $G(\cdot)$ in Remark 2, the normalized near-field interference generally decreases as $\beta_{2}$ increases. From (46), we see that the rotation angle influences the value of the sine function that attains its maximum value of one at $\frac{\pi}{2}$. Therefore, we have
@@
    $\phi^{*}=\begin{cases}\min(\theta-\frac{\pi}{2},\phi_{\rm max}),&{\rm if}\leavevmode\nobreak\ \theta>\frac{\pi}{2},\\ \max(\theta-\frac{\pi}{2},\phi_{\rm min}),&{\rm if}\leavevmode\nobreak\ \theta<\frac{\pi}{2}.\end{cases}$ (47)
@@ Appendix C Proof of Lemma 2
since $\sin{\theta_{k}}-\sin{\theta_{i}}>0$. Thus, $\Delta_{1}(-\epsilon)>0$, implying $\beta_{1}>\bar{\beta}_{1}$.
For $\theta_{i}>\theta_{k}$, the signs of $\sin{(\theta_{k}-\theta_{i})}$ and $\cos{\theta_{i}}-\cos{\theta_{k}}$ reverse, but the analysis holds analogously by adjusting the signs in $\Delta_{1}(\phi)$ and $\Delta_{2}(\phi)$. The case $\theta_{i},\theta_{k}\in(\frac{\pi}{2},\pi)$ is similar, leveraging the symmetry of trigonometric functions. Thus, the inequalities hold for all cases, which completes the proof.

### v4

@@ III-A Closed-form Expressions for Both Interference Terms
A closed-form approximation for the rotation-dependent normalized near-field inter-user interference in (15) is provided in Theorem 1.
@@
The rotation-dependent normalized near-field inter-user interference $\rho_{\rm NN}(\phi,\theta_{k},r_{k},\theta_{i},r_{i})$ in (15) can be approximated as:
@@
Moreover, $\widehat{C}(\beta_{1},\beta_{2})={C}(\beta_{1}+\beta_{2})-{C}(\beta_{1}-\beta_{2})$ and $\widehat{S}(\beta_{1},\beta_{2})={S}(\beta_{1}+\beta_{2})-{S}(\beta_{1}-\beta_{2})$, where $C(\cdot)$ and $S(\cdot)$ are the Fresnel integrals, defined as $C(x)=\int^{x}_{0}\cos(\frac{\pi}{2}t^{2})\text{d}t$ and $S(x)=\int^{x}_{0}\sin(\frac{\pi}{2}t^{2})\text{d}t$, respectively.
@@
Next, we provide a closed-form approximation for the rotation-dependent normalized mixed-field inter-user interference.
@@
The rotation-dependent normalized mixed-field inter-user interference $\rho_{\rm NF}(\phi,\theta_{k},r_{k},\psi_{m})$ in (16) can be approximated as
@@
It is worth highlighting that Theorem 1 generalizes Theorem 2, reducing to it as $r_{k}$ (or $r_{i}$) approaches infinity, i.e., $r_{k}({\rm or}\penalty 10000\ r_{i})\rightarrow\infty$ . Moreover, since both closed-form expressions accounting for the near-field and mixed-field interference are derived capitalizing on the Fresnel integrals, they share the same properties of function $G(\cdot)$, which facilitate the subsequent interference analysis, as detailed below.
@@
      $\rho_{\rm NN}(0,\theta_{k},r_{k},\theta_{i},r_{i})=G(\bar{\beta}_{1},\bar{\beta}_{2})$ [12], where
@@
      $\rho_{\rm NF}(0,\theta_{k},r_{k},\psi_{m})=G(\tilde{\beta}_{1},\tilde{\beta}_{2})$ [9], where
@@
Theorems 1 and 2 reveal an important insight: compared to conventional fixed-antenna systems, the rotation angle $\phi$ introduces an additional DoF for controlling the values of $\beta_{1}$ and $\beta_{2}$, and consequently, the interference levels measured by $G(\cdot)$. This manifests that the added flexibility enables the suppression of both near-field and mixed-field interference through strategic adjustment of the rotation angle $\phi$. In the subsequent subsections, we analytically demonstrate how the rotation angle $\phi$ affects the near-field and mixed-field interference, respectively, and hence the overall system performance.
@@ III-B Near-field Inter-user Interference
For two near-field users located at the same angle, i.e., $\theta_{k}=\theta_{i}=\theta$, but different distances, i.e., $r_{k}\neq r_{i}$, the rotation-dependent normalized near-field interference in (17) depends solely on $\beta_{2}$ in (1), such that $\rho_{\rm NN}\approx G(0,\beta_{2})$. The effect of array rotation on this interference can be divided into the following two cases:
@@
    $\phi^{*}=\begin{cases}\min(\theta-\frac{\pi}{2},\phi_{\rm max}),&{\rm if}\penalty 10000\ \theta>\frac{\pi}{2},\\ \max(\theta-\frac{\pi}{2},\phi_{\rm min}),&{\rm if}\penalty 10000\ \theta<\frac{\pi}{2}.\end{cases}$ (24)
@@
To demonstrate the benefits of RAs in mitigating the near-field inter-user interference, we first validate the accuracy of the proposed approximation in Theorem 1 and the results derived in Lemma 1. This corroboration is presented in Fig. 5, where the Fresnel-based approximation is shown to closely match the exact normalized interference given in (15). Moreover, the derived rotation angles align closely with the numerical values across various spatial-angle settings. Next, we investigate the effects of the rotation angle on the achievable sum-rate with respect to the spatial angle and distance of the near-field user. This analysis is detailed in Figs. 55, where the system parameters are set to $N=129$, $f=24$ GHz, and $K=2$ (i.e., near-field users NU1 and NU2). We see that RA-enabled systems consistently outperform their fixed-antenna counterparts across all tested angles and distances. We explore the effect of rotation on the achievable sum-rate from both angular and distance perspectives as follows:
@@
  • Effect of angle given the same distance: Fig. 5 illustrates that when two near-field users are located at the same distance, the performance gain from RAs depends heavily on their angular separation. Notably, significant gains are observed when the angular separation is small, where interference is inherently strong, and RAs can effectively reduce it by adjusting the array’s orientation. As the angular separation increases, both RA-enabled and fixed-antenna systems benefit from reduced interference due to the increased orthogonality of the near-field and far-field steering vectors.
@@ III-C Mixed-field Inter-user Interference
When the far-field user and the near-field user are at the same angle, i.e., $\psi=\theta$, the rotation-dependent normalized mixed-field interference in (19) depends only on $\beta_{2}$ in (2), such that $\rho_{\rm NF}\approx G(0,\beta_{2})$. The effect of rotation on this interference can be characterized as follows:
@@
    $\phi^{*}=\begin{cases}\min(\psi-\frac{\pi}{2},\phi_{\rm max}),&{\rm if}\penalty 10000\ \psi>\frac{\pi}{2},\\ \max(\psi-\frac{\pi}{2},\phi_{\rm min}),&{\rm if}\penalty 10000\ \psi<\frac{\pi}{2}.\end{cases}$ (27)
@@
  • When the near-field and far-field users share the same spatial angle but are not at the boresight direction (i.e., $\theta=\psi\neq\frac{\pi}{2}$), performance gains from array rotation depend on the spatial angle, which diminish when the angle approaches the boresight direction from either side.
@@ IV-A Inner Problem: Power Allocation Optimization
where the gradient of $B_{1}$ in terms of $\{P_{{\rm N},k}\}$ is given by (32), as shown on the bottom of this page. The resultant problem is convex with respect to $\{P_{{\rm N},k}\}$, and thus can be efficiently solved by standard convex solvers such as CVX.
@@
IV-B Outer Problem: Rotation Angle Optimization
@@
$f_{\rm fit}(\bm{\phi}^{(t)}_{s})=\sum_{k\in\mathcal{K}}\penalty 10000\ R_{k}(\bm{\phi}^{(t)}_{s})-\xi\mathbb{P}(\bm{\phi}^{(t)}_{s})$, (39)
@@ V Numerical Results
  • Rotatable antenna with ZF digital beamforming (RA+ZF): The rotation angle is optimized using the proposed algorithm, while ZF beamforming is applied.
@@ V-A Sum-rate Versus Maximum Transmit Power
In Fig. 9, we illustrate the effect of the transmit power on the sum-rate performance of various schemes. First, it is observed that the proposed scheme closely matches the performance of its combination with ZF digital beamforming and consistently outperforms the benchmark schemes, highlighting the effectiveness of RAs in enhancing the mixed-field communication performance. Second, compared to the RA+EPA scheme, which optimizes rotation alone, joint optimization of rotation and power allocation yields a higher sum-rate, underscoring the efficacy of the integrated approach. Third, subarray-wise rotation delivers significant performance gains over the array-wise rotation, owing to its superior interference mitigation capabilities enabled by greater design flexibility. Additionally, one interesting observation is that the FA+ZF scheme underperforms the proposed method across all transmit power levels. This is because, while FA+ZF effectively nullifies the near-field inter-user interference, it simultaneously suffers from the complex mixed-field interference caused by far-field users, resulting in notable performance degradation, especially when the mixed-field interference is pronounced.
@@ VI Conclusion
In this paper, we proposed leveraging RAs to mitigate near-field and mixed-field interference by exploiting the spatial DoF enabled by RA array rotation. For the special case where all subarrays share a uniform rotation angle, we derived the closed-form expressions for the rotation-dependent near-field interference and the rotation-dependent mixed-field interference using the Fresnel integrals. In particular, we confirmed that array rotation effectively suppresses complex near-field and mixed-field interference, thus achieving substantial performance improvements in mixed-field communication scenarios. For the general case with subarray-wise rotation, we have developed a double-layer algorithm, where the inner layer optimizes power allocation via SCA, and the outer layer determines subarray rotation angles via PSO. Numerical results demonstrated that the RA-enabled system substantially outperforms conventional fixed-antenna configurations and underscore the superior effectiveness of our proposed joint design compared to benchmark schemes.
Appendix A
@@
Based on the properties of the function $G(\cdot)$ in Remark 2, the normalized near-field interference generally decreases as $\beta_{2}$ increases. From (46), we see that the rotation angle influences the value of the sine function that attains its maximum value of one at $\frac{\pi}{2}$. Therefore, we have
@@
    $\phi^{*}=\begin{cases}\min(\theta-\frac{\pi}{2},\phi_{\rm max}),&{\rm if}\penalty 10000\ \theta>\frac{\pi}{2},\\ \max(\theta-\frac{\pi}{2},\phi_{\rm min}),&{\rm if}\penalty 10000\ \theta<\frac{\pi}{2}.\end{cases}$ (47)
@@
since $\sin{\theta_{k}}-\sin{\theta_{i}}>0$. Thus, $\Delta_{1}(-\epsilon)>0$, implying $\beta_{1}>\bar{\beta}_{1}$. For $\theta_{i}>\theta_{k}$, the signs of $\sin{(\theta_{k}-\theta_{i})}$ and $\cos{\theta_{i}}-\cos{\theta_{k}}$ reverse, but the analysis holds analogously by adjusting the signs in $\Delta_{1}(\phi)$ and $\Delta_{2}(\phi)$. The case $\theta_{i},\theta_{k}\in(\frac{\pi}{2},\pi)$ is similar, leveraging the symmetry of trigonometric functions. Thus, the inequalities hold for all cases, which completes the proof.
Appendix D Acknowledgment
The authors would like to thank the developers of Qwen, an advanced large language model, for providing valuable assistance during the language polish process.